# SWAP-less Implementation of Quantum Algorithms

### v1

@@ IV.1 Quantum Fourier Transform
where the subscript $jk$ indicates the qubits it acts on. We are in particular interested in implementing the QFT on a device with linear nearest-neighbor (LNN) connectivity, a problem that was already studied in several works [35, 36, 37, 38, 8, 7], requiring $23n^{2}/$ 2 CNOT-gates in leading order [8] or introducing a quadratic circuit depth [7]. With the above scheme we implement the QFT on an LNN architecture with ${n^{2}-1}$ CNOT gates and a total circuit depth of ${5n-3}$, which outperforms all other current approaches to the best of our knowledge.
Our approach starts with a set of qubits holding the initial quantum state and then applies the QFT in the parity architecture as presented in Ref. [39] by iterating through different spanning lines. For that we apply sequences of CNOT gates as depicted in Fig. 3, which illustrates the corresponding circuit for ${n=5}$ qubits. The logical CPhase gates in the QFT circuit are decomposed into three $R_{Z}$ rotations on base- and parity qubits according to
@@
When using previous approaches for realizing the QFT, it may be necessary to reverse the qubit order afterwards if the QFT is not at the end of the total quantum circuit [26]. This requires $3\lfloor\frac{n}{2}\rfloor$ CNOT gates and a single time step on an all-to-all connected device and $23n(n-1)/$ 2 CNOT gates and $3n$ time steps on an architecture with LNN connectivity, where the factor of 3 originates from decomposing SWAP gates into CNOT gates. As our approach intrinsically reverses the order of qubits throughout the algorithm execution (see Fig. 3), we eliminate another quadratic amount of CNOT gates on top of the savings in the actual QFT circuit, when taking the correct ordering of qubits into account.
@@
{tblr}
colspec=—l—cccc—, rowsep=1.5pt, colsep=2.5pt, hline1,2,5,8 Resource & Parity Ref. [7] Ref. [8] all-to-all [8]
@@
CNOT depth $44n-$ 4 $n^{2}+n-4$ $96n-$ 9 $64n-$ 6
SQ depth $n+1$ $12n+$ 1 $44n-$ 4 $44n-$ 4
total depth $35n-$ 3 $n^{2}+3n-3$ $1310n-$ 13 $108n-$ 10
Table 1: Comparison of the required resources for the QFT on $n$ qubits in a linear chain using different approaches, and additionally to an all-to-all connected system. The gate set used comprises CNOT gates as well as single-qubit (SQ) $H$, $R_{x}$ and $R_{z}$ gates. Note that, if SQ gates are not executed in parallel with CNOT-gates in our approach, there are $12n-$ 1 steps with solely SQ-gates. In that case, the total depth increases to $56n-$ 5.
@@ IV.2 The Quantum Approximate Optimization Algorithm
|$\ket{\psi(\bm{\beta},\bm{\gamma})}=\prod_{j=1}^{p}U_{X}(\beta_{j})U_{P}(\gamma_{j})\ket{+}^{\otimes n},$ (17)
@@
$E(\bm{\beta},\bm{\gamma})=\braket{\psi(\bm{\beta},\bm{\gamma})}{H_{\text{P}}}{\psi(\bm{\beta},\bm{\gamma})}.$ (18)
@@ IV.2.1 The QAOA on a linear chain
To perform the QAOA on a chain with linear nearest-neighbor (LNN) connectivity, typically SWAP networks [42, 43, 12, 44, 45] are used to ensure connectivity between all interacting qubits throughout the algorithm. If all-to-all connectivity is required, the optimal SWAP network results in an overhead of $n^{2}/2$ CNOT gates in addition to the CNOT gates necessary to implement the interaction terms, resulting in a total number of required CNOT gates of $23n^{2}/$ 2 in leading order [9], while maintaining a linear circuit depth.
@@
In the temporal parity encoding, $U_{P}$ is implemented by first sequentially transforming an initial zig-zag shaped spanning line, as illustrated in Fig. 2(a) by the qubits and lines with the thick borders, to ${\lceil(n-1)/2\rceil}$ other similarly shaped spanning lines. For that, we use the circuit design in Fig. 2(b). This corresponds to covering all the labels present in half a period of the extended LHZ-code  [^2] For an even number of qubits, the last spanning line only needs to differ from the previous one in half of the qubits and its creation only requires ${n-1}$ CNOT gates in 2 time steps. Therefore, the final numbers are independent of the parity of $n$.. By construction, these spanning lines cover all necessary (and, in this case, all possible) labels ${\{j,k\}}$, which then allows for the execution of the desired logical operations $\exp(-i\gamma J_{jk}\bar{Z}_{j}\bar{Z}_{k})$ as physical single-qubit rotations $R_{Z}$ according to Eq. (9). This results in a CNOT depth of ${2n-2}$ and a CNOT gate count of ${(n-1)^{2}}$ (cf. Sec. III.2).
@@
Figure 4: (a) The quantum circuit to implement the problem unitary $U_{P}$ of the QAOA with parameter $\gamma$ by means of the time-encoded parity layout, where the angles are defined as ${\tilde{J}_{ij}=\gamma{J}_{ij}}$ and ${\tilde{h}_{i}=\gamma{h}_{i}}$. The circuit propagates through the different spanning lines as indicated by the changing labels and logical lines (colored lines in the figure). (b) The quantum circuit to implement the mixer unitary $U_{X}$ with parameter $\beta$ of the QAOA. This circuit does not change the spanning line and requires ${2(n-2)}$ CNOT gates during $4$ time steps and $n$ single-qubit gates in $2$ time steps.
@@
The unitary $U_{X}$ can be applied using logical $R_{X}(\beta_{j})$ gates along the logical lines as defined in Eq. (10). The logical lines at a given time always span at most two neighboring qubits and a spanning line always has at least one localized logical $X$-operator (i.e., logical line of length 1) on a boundary. Thus, applying $U_{X}$ requires a constant two-qubit gate depth of 44{4}4 and at most ${2n-2}$ CNOT gates. The circuit diagram for implementing $U_{X}$ on an exemplary spanning line is given in Fig. 4(b).
@@
In total, our approach exhibits a CNOT-gate count of $n^{2}-1$ and with that outperforms the SWAP network by a factor of $3/2$ in leading order. Furthermore, it results in a shorter circuit depth while the number of required single-qubit rotations remains unchanged. The numbers for gate counts and circuit depth in terms of the universal gate set $\{R_{z},R_{x},H,\text{CNOT}\}$ are given in Tab. IV.2.1. These numbers also hold for the trotterized time evolution of the transverse-field Ising Hamiltonian, as there is a direct relation to the QAOA.
Note that we have not taken into account the initialization of the logical state |$\ket{+}^{\otimes n}$ in our discussion. However, this state corresponds to all qubits in the spanning line being in the state $\ket{+}$. Therefore, it can be prepared by simply applying physical Hadamard gates on all physical qubits prepared in the $|0\rangle$ state.
@@
Remarkably, also on two-dimensional grids, no strategies have been found that outperform the linear SWAP strategy [13]. Consequently, our strategy reduces the resource requirements for QAOA not only on a LNN chain, but also on devices with 2D connectivity.
{tblr}
colspec=l—l—ccc—, rowsep=1.5pt, colsep=2.5pt, hline2,6,12, hline1 = 2-Zsolid, vline1 = 2-Zsolid, cell21=r=4c, cell61=r=6c, cell121=r=6c, & Resource Linear Parity Ladder-Parity Ref. [9]
$U_{P}$CNOTs ${n^{2}}-2n+1$ ${n^{2}}-n+1$ ${\frac{3}{2}n^{2}}-\frac{5}{2}n+1$
CNOT depth $22n-$ 2 $2\lceil\frac{n}{2}\rceil$ $23n-$ 2
#SQ gates $n(n+1)/2$ $n(n+1)/2$ $n(n-1)/2$
@@
$U_{X}$CNOTs $22n-$ 2 $22n-$ 2 00
CNOT depth $4$ $4$ 00
@@
$U_{X}U_{P}$CNOTs $n^{2}-1$ $n^{2}+n-1$ $\frac{3}{2}n^{2}-\frac{5}{2}n+1$
CNOT depth $22n+$ 2 $42\lceil\frac{n}{2}\rceil+$ 4 $23n-$ 2
#SQ gates $\frac{1}{2}n(n+1)$ $\frac{1}{2}n(n+2)$ $\frac{1}{2}n(n+1)$
@@
Table 2: Measurement, CNOT and single qubit (SQ) depth and gate counts for the building blocks of the QAOA for an all-to-all connected problem graph of size $n$ on a device with LNN or ladder connectivity. The used gate set comprises CNOT gates as well as single-qubit $H$, $R_{X}$ and $R_{Z}$ gates. Note that in Ref. [9] no local field terms are included. Since our presented protocol includes local field terms, the same circuit could also be used to implement the QAOA for a problem of size ${n+1}$ without local field terms.
@@ IV.2.2 The QAOA on a ladder
If $2n$ physical qubits are available and arranged in a ladder geometry (i.e., a $2\times n$ grid with nearest neighbor connectivity), then both rails of the ladder can be used to progress to new spanning lines in parallel, by using the circuit in Fig. 4(a) independently on each rail. This allows us to reach all required qubit labels in half the circuit depth compared to the approach presented in Sec. IV.2.1. We begin by assuming an arbitrary spanning line on the first rail and empty qubits on the second. The second rail is initialized by “copying” the labels of the spanning line from one rail to the other rail, using the CNOT circuit shown in Fig. 5(b). This circuit applies $n$ CNOT gates in parallel, where each CNOT gate is controlled by one qubit in the first rail and targets onto its neighboring empty qubit on the second rail of the ladder. As a result, we obtain two spanning lines with identical labels.
@@
Then, we successively progress the two spanning lines in opposite directions by applying the quantum circuit from Fig. 4(a) to each rail, as depicted in Fig. 5(a). For one of the spanning lines, the targets and controls of each CNOT gate are switched, which reverses the direction of progression for this spanning line. Note that for some $n$, it can be beneficial to initially progress each rail by half a step [see transition from $\tau_{0}$ to $\tau_{1}$ in Fig. 5(a)] before applying the first $Z$ rotations in order to reach more parities in that round already. This approach requires $n$ CNOT gates for the “copying” and $(n-1)^{2}$ CNOT gates for the progression, with a total circuit depth of $2\lceil\frac{n}{2}\rceil$ until all parities have been reached at some step.
@@
After applying all interactions in $U_{P}$, the logical lines are delocalized as depicted in Fig. 5(c). To relocalize them such that the logical lines occur along two neighboring qubits along on a single rail, we measure all qubits in one rail in the $X$-basis and apply conditional $Z$ gates to the qubits on the other rail. Such a protocol has the same effect as the decoding CNOT sequence required to remove all the labels from the first rail. We refer the reader to Ref. [25] for details on the classical corrections necessary and to Ref. [47] for the relation to measurement-based quantum computing. This step effectively decodes one rail of the ladder, such that no redundant information remains. The decoding process adds $n$ qubit measurements and $2n$ single-qubit rotations to the resource requirements of $U_{X}$. After decoding, $U_{X}$ can be applied on the single remaining spanning line using the same circuit components and therefore the same resource requirements as in Sec. IV.2.1. With that, one QAOA round is finished and we can, if desired, continue by copying the labels of the current spanning line to the second ladder rail again and starting a new round from there. The overall resource requirements for this method are given in Tab. IV.2.1.
@@
Figure 5: (a) The effective evolution of two spanning lines in opposite directions for applying $U_{P}$ on a ladder architecture for ${n=6}$. The logical lines with dotted borders indicate the initial spanning line ($\tau_{0})$. (b) The initialization of a second spanning line from a first spanning line by using empty qubits and $n$ CNOT gates. (c) The two spanning lines after implementing $U_{P}$, where the logical lines are delocalized. In order to allow easier implementation of $U_{X}$, the encoding is reduced to a single spanning line again, by decoding the second spanning line via $X$-basis measurements and conditional $Z$-gates.
@@ V The use of other entangling gates
For devices where the CNOT gate is not a native gate, one can reformulate the above-presented circuits by expressing a CNOT gate as a controlled-Z gate or a controlled-phase gate preceded and followed by a Hadamard gate or a Y-rotation on the target qubit. The resource counts from Tables 1 and IV.2.1 still hold up in terms of entangling-gate counts and depths.
@@ VI Conclusion and Outlook
In this work, we have formalized the tracking of parity labels and demonstrated its utility in quantum circuit design. For the QFT on a linear chain, we find lower resource requirements in terms of two-qubit gate count and depth for the gate set $\{R_{z},R_{x},H,\text{CNOT}\}$ with respect to various other published works and even the same resource requirements as for an all-to-all connected system in leading order. Our scheme reduces the two-qubit gate count and depth of the QAOA on a LNN architecture by a factor of $2/3$ in leading order compared to previously published methods. Moreover, our implementation of the QAOA on a ladder architecture, reducing circuit depth by a factor of two without increasing the gate count in leading order, trades off qubit count for circuit depth. These advancements highlight a promising approach for implementing near-term quantum algorithms on NISQ devices, where minimizing gate count and circuit depth is crucial.
Further research may involve the development of a compilation tool to leverage these findings for algorithms that do not require all-to-all connectivity. Additionally, it would be valuable to investigate the actual improvements in algorithmic performance and the reliability of quantum circuit outputs on real quantum devices. Building on our analytical evaluation of the necessary resources in terms of CNOT-gate counts and depth, we anticipate significant improvements not only within the theoretical framework but also in a practical experimental setting, e.g., when subjected to quantum noise. Finally, we note that the tracking approach can be generalized to include other Clifford gates, related to a concept known as Clifford tableaus [49].

### v2

@@ IV.1 Quantum Fourier Transform
where the subscript $jk$ indicates the qubits it acts on. We are in particular interested in implementing the QFT on a device with linear nearest-neighbor (LNN) connectivity, a problem that was already studied in several works [36, 37, 38, 39, 8, 7], requiring $3n^{2}/2$ CNOT gates in leading order [8] or introducing a quadratic circuit depth [7]. With the above scheme we implement the QFT on an LNN architecture with $n^{2}-1$} CNOT gates and a total circuit depth of 5n3{5n-3}, which outperforms all other current approaches to the best of our knowledge.
Our approach starts with a set of qubits holding the initial quantum state and then applies the QFT in the parity architecture as presented in Ref. [40] by iterating through different spanning lines. For that we apply sequences of CNOT gates as depicted in Fig. 3, which illustrates the corresponding circuit for n=5{n=5} qubits. The logical CPhase gates in the QFT circuit are decomposed into three $R_{Z}$ rotations on base- and parity qubits according to
@@
When using previous approaches for realizing the QFT, it may be necessary to reverse the qubit order afterwards if the QFT is not at the end of the total quantum circuit [26]. This requires $3\lfloor\frac{n}{2}\rfloor$ CNOT gates and a single time step on an all-to-all connected device and 3n(n1)/23n(n-1)/2 CNOT gates and $3n$ time steps on an architecture with LNN connectivity, where the factor of 3 originates from decomposing SWAP gates into CNOT gates. As our approach intrinsically reverses the order of qubits throughout the algorithm execution (see Fig. 3), we eliminate another quadratic amount of CNOT gates on top of the savings in the actual QFT circuit, when taking the correct ordering of qubits into account.
@@
Resource Parity Ref. [7] Ref. [8] all-to-all [8]
@@
CNOT depth 4n44n-4 $n^{2}+n-4$ 6n96n-9 4n64n-6
SQ depth $n+1$ $2n+1$ 4n44n-4 4n44n-4
total depth 5n35n-3 $n^{2}+3n-3$ 10n1310n-13 8n108n-10
Table 1: Comparison of the required resources for the QFT on $n$ qubits in a linear chain using different approaches, and additionally to an all-to-all connected system. The gate set used comprises CNOT gates as well as single-qubit (SQ) $H$, $R_{X}$ and $R_{Z}$ gates. Note that, if SQ gates are not executed in parallel with CNOT gates in our approach, there are 2n12n-1 steps with solely SQ gates. In that case, the total depth increases to 6n56n-5.
@@ IV.2 The Quantum Approximate Optimization Algorithm
[…]
@@
$E(\bm{\beta},\bm{\gamma})=\braket{\psi(\bm{\beta},\bm{\gamma})|H_{\text{P}}|\psi(\bm{\beta},\bm{\gamma})}.$ (18)
@@ IV.2.1 The QAOA on a linear chain
To perform the QAOA on a chain with linear nearest-neighbor (LNN) connectivity, typically SWAP networks [43, 44, 12, 45, 46] are used to ensure connectivity between all interacting qubits throughout the algorithm. If all-to-all connectivity is required, the optimal SWAP network results in an overhead of $n^{2}/2$ CNOT gates in addition to the CNOT gates necessary to implement the interaction terms, resulting in a total number of required CNOT gates of $3n^{2}/2$ in leading order [9], while maintaining a linear circuit depth.
@@
In the temporal parity encoding, $U_{P}$ is implemented by sequentially transforming an initial zig-zag shaped spanning line to $\lceil(n-1)/2\rceil$} other spanning lines and applying single-qubit $Z$ rotations in between each transformation. The transformation between two zig-zag shaped spanning lines is illustrated in Fig. 2(a) by the qubits and lines with the thick borders and it is implemented via the circuit design in Fig. 2(b). This method corresponds to covering all the labels present in half a period of the extended LHZ code  [^2] For an even number of qubits, the last spanning line only needs to differ from the previous one in half of the qubits and its creation only requires n1{n-1} CNOT gates in 2 time steps. Therefore, the final numbers are independent of the parity of $n$.. By construction, these spanning lines cover all necessary (and, in this case, all possible) labels $\{j,k\}$}, which allows for the execution of the desired logical operations $\exp(-i\gamma J_{jk}\bar{Z}_{j}\bar{Z}_{k})$ as physical single-qubit rotations $R_{Z}$ according to Eq. (9). This implementation of $U_{P}$ is shown in Fig. 4(a) and results in a CNOT depth of 2n2{2n-2} and a CNOT gate count of $(n-1)^{2}$} (cf. Sec. III.2).
@@
Figure 4: (a) The quantum circuit to implement the problem unitary $U_{P}$ of the QAOA with parameter $\gamma$ by means of the time-encoded parity layout, where the angles are defined as $\tilde{J}_{ij}=\gamma{J}_{ij}$} and $\tilde{h}_{i}=\gamma{h}_{i}$}. The circuit propagates through the different spanning lines as indicated by the changing labels and logical lines (colored lines in the figure). (b) The quantum circuit to implement the mixer unitary $U_{X}$ with parameter $\beta$ of the QAOA. This circuit does not change the spanning line and requires 2(n2){2(n-2)} CNOT gates during $4$ time steps and $n$ single-qubit gates in 22 time steps.
@@
The unitary $U_{X}$ can be applied using logical $R_{X}(\beta_{j})$ gates along the logical lines as defined in Eq. (10). The logical lines at a given time always span at most two neighboring qubits and a spanning line always has at least one localized logical $X$ operator (i.e., logical line of length 1) on a boundary. Thus, applying $U_{X}$ requires a constant two-qubit gate depth of 4{4} and at most 2n2{2n-2} CNOT gates. The circuit diagram for implementing $U_{X}$ on an exemplary spanning line is given in Fig. 4(b).
@@
In total, our approach exhibits a CNOT gate count of $n^{2}-1$ and with that outperforms the SWAP network by a factor of $3/2$ in leading order. Furthermore, it results in a shorter circuit depth while the number of required single-qubit rotations remains unchanged. The numbers for gate counts and circuit depth in terms of the universal gate set $\{R_{Z},R_{X},H,\text{CNOT}\}$ are given in Tab. 2. These numbers also hold for the trotterized time evolution of the transverse-field Ising Hamiltonian, as there is a direct relation to the QAOA.
[…]
@@
Remarkably, also on two-dimensional grids, no strategies have been found that outperform the linear SWAP strategy [13]. Consequently, our strategy reduces the resource requirements for QAOA not only on an LNN chain, but also on devices with 2D connectivity.
Resource Linear Parity Ladder-Parity Ref. [9]
$U_{P}$ #CNOTs ${n^{2}}-2n+1$ ${n^{2}}-n+1$ ${\frac{3}{2}n^{2}}-\frac{5}{2}n+1$
CNOT depth 2n22n-2 $2\lceil\frac{n}{2}\rceil$ 3n23n-2
#SQ gates $n(n+1)/2$ $n(n+1)/2$ n(n1)/2n(n-1)/2
@@
$U_{X}$ #CNOTs 2n22n-2 2n22n-2 0
CNOT depth 44 44 0
@@
$U_{X}U_{P}$ #CNOTs $n^{2}-1$ $n^{2}+n-1$ $\frac{3}{2}n^{2}-\frac{5}{2}n+1$
CNOT depth $2n+2$ $2\lceil\frac{n}{2}\rceil+4$ 3n23n-2
#SQ gates $n(n+3)/2$ $n(n+5)/2$ $n(n+1)/2$
@@
Table 2: Measurement, CNOT and single-qubit (SQ) depth and gate counts for the building blocks of the QAOA for an all-to-all connected problem graph of size $n$ on a device with LNN or ladder connectivity. The used gate set comprises CNOT gates as well as single-qubit $H$, $R_{X}$ and $R_{Z}$ gates. Note that in Ref. [9] no local field terms are included. Since our presented protocol includes local field terms, the same circuit could also be used to implement the QAOA for a problem of size n+1{n+1} without local field terms.
@@ IV.2.2 The QAOA on a ladder
If $2n$ physical qubits are available and arranged in a ladder geometry (i.e., a $2\times n$ grid with nearest-neighbor connectivity), then both rails of the ladder can be used to progress to new spanning lines in parallel, by using the circuit in Fig. 4(a) independently on each rail. This allows us to reach all required qubit labels in half the circuit depth compared to the approach presented in Sec. IV.2.1. The gate depth of the full QAOA circuit is then reduced to that of an all-to-all connected system [48], at the cost of doubling the qubit number and gate count, and introducing a linear number of single-qubit mid-circuit measurements. For the use of measurements in parity code deformations we refer the reader to Refs. [25, 49].
To allow for parallel evolution of two logical lines, we begin by assuming an arbitrary spanning line on the first rail and empty qubits on the second. The second rail is initialized by “copying” the labels of the spanning line from one rail to the other rail, using the CNOT circuit shown in Fig. 5(b). This circuit applies $n$ CNOT gates in parallel, where each CNOT gate is controlled by one qubit in the first rail and targets onto its neighboring empty qubit on the second rail of the ladder. As a result, we obtain two spanning lines with identical labels.
@@
Then, we successively progress the two spanning lines in opposite directions by applying the quantum circuit from Fig. 4(a) to each rail, as depicted in Fig. 5(a). For one of the spanning lines, the targets and controls of each CNOT gate are switched, which reverses the direction of progression for this spanning line. Note that for some $n$, it can be beneficial to initially progress each rail by half a step [see transition from $\tau_{0}$ to $\tau_{1}$ in Fig. 5(a)] before applying the first $Z$ rotations in order to reach more parities in that round already. This approach requires $n$ CNOT gates for the “copying” and $(n-1)^{2}$ CNOT gates for the progression, with a total circuit depth of $2\lceil\frac{n}{2}\rceil$ until all parities have been reached at some time step.
@@
After applying all interactions in $U_{P}$, the logical lines are delocalized as depicted in Fig. 5(c). To relocalize them such that the logical lines occur along two neighboring qubits on a single rail, we measure all qubits in one rail in the $X$ basis and apply conditional-$Z$ gates to the qubits on the other rail. Such a protocol has the same effect as the decoding CNOT sequence required to remove all the labels from the first rail (see [25] for the necessary classical corrections). This step effectively decodes one rail of the ladder, such that no redundant information remains. The decoding process adds $n$ qubit measurements and $2n$ single-qubit rotations to the resource requirements of $U_{X}$. After decoding, $U_{X}$ can be applied on the single remaining spanning line using the same circuit components and therefore the same resource requirements as in Sec. IV.2.1. With that, one QAOA round is finished and we can, if desired, continue by copying the labels of the current spanning line to the second ladder rail again and starting a new round from there. The overall resource requirements for this method are given in Tab. 2.
@@
Figure 5: (a) The effective evolution of two spanning lines in opposite directions for applying $U_{P}$ on a ladder architecture for n=6{n=6}. The logical lines with dotted borders indicate the initial spanning line ($\tau_{0})$. (b) The initialization of a second spanning line from a first spanning line by using empty qubits and $n$ CNOT gates. (c) The two spanning lines after implementing $U_{P}$, where the logical lines are delocalized. In order to allow easier implementation of $U_{X}$, the encoding is reduced to a single spanning line again, by decoding the second spanning line via $X$-basis measurements and conditional $Z$ gates.
@@ V The use of other entangling gates
For devices where the CNOT gate is not a native gate, one can reformulate the above-presented circuits by expressing a CNOT gate as a controlled-$Z$ gate or a controlled-phase gate preceded and followed by a Hadamard gate or a $Y$ rotation on the target qubit. The resource counts from Tables 1 and 2 still hold up in terms of entangling-gate counts and depths.
@@ VI Conclusion and Outlook
In this work, we have formalized the tracking of parity labels and demonstrated its utility in quantum circuit design. For the QFT on a linear chain, we find lower resource requirements in terms of two-qubit gate count and depth for the gate set $\{R_{Z},R_{X},H,\text{CNOT}\}$ with respect to various other published works and even the same resource requirements as for an all-to-all connected system in leading order. Our scheme reduces the two-qubit gate count and depth of the QAOA on an LNN architecture by a factor of $2/3$ in leading order compared to previously published methods. Moreover, our implementation of the QAOA on a ladder architecture, reducing circuit depth by a factor of two without increasing the gate count in leading order, trades off qubit count and mid-circuit measurements for circuit depth. These advancements highlight a promising approach for implementing near-term quantum algorithms on NISQ devices, where minimizing gate count and circuit depth is crucial. This approach has also been investigated beyond the LNN architecture, where it can exploit the additional connectivity to further reduce two-qubit gate counts [48]. Further research may involve the development of a compilation tool to leverage these findings for algorithms that do not require all-to-all connectivity. Additionally, it would be valuable to investigate the actual improvements in algorithmic performance and the reliability of quantum circuit outputs on real quantum devices. Building on our analytical evaluation of the necessary resources in terms of CNOT gate counts and depth, we anticipate significant improvements not only within the theoretical framework but also in a practical experimental setting, e.g., when subjected to quantum noise. Finally, we note that the tracking approach can be generalized to include other Clifford gates and Pauli rotations, related to a concept known as Clifford tableaus [51] and flow tableaus [35].